( 81 ) - 7/4 _ ( 81 ) -3/4 divide ( 81 ) 5/4

 ( 81 ) - 7/4      _   ( 81 ) -3/4 
 
         divide  
 
(  81 ) 5/4

Grade:9

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given
\frac{81^{-7/4}-81^{-3/4}}{81^{5/4}}
By shifting denominator to numerator by using property am/an=am-nwe get
=(81)(-7/4)-(5/4)-(81)(-3/4)-(5/4)
=(81)(-12/4)-(81)(-8/4)
=(34)(-12/4)-(34)(-8/4)
=3-12-3-8
=(1/312)-(1/38)
=(312-38)/312
=0.9876